# Indoor Light and Heat Estimation from a Single Panorama

### v1

@@ Abstract
This paper presents a novel application for directly estimating indoor light and heat maps from captured indoor-outdoor High Dynamic Range (HDR) panoramas. In our image-based rendering method, the indoor panorama is used to estimate the 3D room layout, while the corresponding outdoor panorama serves as an environment map to infer spatially-varying light and material properties. We establish a connection between indoor light transport and heat transport and implement transient heat simulation to generate indoor heat panoramas. The sensitivity analysis of various thermal parameters is conducted, and the resulting heat maps are compared with the images captured by the thermal camera in real-world scenarios. This digital application enables automatic indoor light and heat estimation without manual inputs and cumbersome field measurements.
keywords:
Indoor Scene; Light Estimation; Heat Estimation; Image-Based Rendering; HDR Photography1 Introduction
Outdoor light enters indoor spaces, shapes the appearance of indoor scenes, and brings solar radiation that generates heat on indoor surfaces. Omnidirectional photography has been used to capture panoramic images of entire indoor scenes. Based on physics-based and data-driven methods (Ji et al. (2023, 2024); Zhi et al. (2022)), a single indoor panorama can be directly used to estimate spatially varying indoor lighting, material properties, and 3D layouts. Estimating indoor light levels allows users to identify optimal areas for activities, while heat estimation helps assess the potential heating and cooling demand for the entire space. However, most thermal imaging relies on infrared thermal cameras, which are expensive, output noisy imaging results, and are limited to capturing 2D perspective images, making it challenging to cover an entire 360∘ indoor scene in high quality. This raises a main research question: Can we estimate indoor lighting and heating properties simultaneously from a single indoor panorama?
@@
The global rendering method proposed by Debevec (2006) offers a High Dynamic Range (HDR) image-based rendering model for relighting virtual objects within a realistic scene context. Virtual sky models (CIE (2003); Perez et al. (1993); Hosek and Wilkie (2012)) have been employed to represent outdoor light in the real world. To reflect instantaneous light estimation for indoor scenes, image-based rendering incorporates realistic sky images as the illumination source. The HDR technique captures the sky appearance, including direct sunlight in outdoor spaces (Stumpfel et al. (2004)). The captured HDR images provide accurate sky luminance distribution and details of the urban context (Inanici (2009, 2010)). In this setting, the indoor scene is accurately rendered using the outdoor sky image at the pixel level. However, so far, the image-based rendering method has not been explored with realistic indoor textures to render photorealistic indoor appearances to estimate absolute light levels and heat maps.
@@
When modeling visible light transport, intrinsic decomposition separates a single image into two layers: a reflectance layer, which captures the material surface’s albedo or color invariant, and a shading layer, which is shaped by light interactions and object geometry. Intrinsic decomposition has been applied in various tasks, including image editing (Bonneel et al. (2017)), removing interreflections (Seitz et al. (2005)), and separating illumination for material recoloring (Carroll et al. (2011)), as well as being applied to multi-view outdoor photos (Duchêne et al. (2015)). To construct indoor-outdoor light transport, Ji et al. (2023, 2024) use an indoor panorama and paired outdoor hemispherical image to virtually render existing scenes with new floor layouts, materials, and illumination conditions. In this study, we estimate the absolute light levels in virtual indoor scenes by modeling light transport using paired indoor and outdoor panoramas. This approach offers an efficient method for accurately simulating indoor illumination conditions.
@@
Heat transport includes the process of heat generated by the source and its exchange between different mediums (Bergman et al. (2011)). Temperature exchange within an object follows the well-known heat transfer physics, and three primary mechanisms include conduction, convection, and radiation. Recent work by Ramanagopal et al. (2024) bridges visible light transport and heat transport in solids and estimates the heat generation from light absorption. Previous studies on heat estimation have primarily focused on small objects under controlled, artificial lighting and captured thermal images in a 2D perspective. In this work, by determining the reflectance properties of indoor materials from captured indoor-outdoor images, the amount of heat absorbed can be estimated.
In this work, we proposed a digital application that exploits paired indoor-outdoor HDR panoramas as input and then estimates indoor lighting and heating properties in panoramic representation (Fig. 2). This work demonstrates that indoor heat generated by light absorption can be estimated by modeling heat transport using a forward rendering process. This work makes the following technical contributions:
An image-based rendering approach using paired indoor-outdoor panoramas to estimate indoor materials, 3D geometry, and lighting for physics-based building analysis.
An application for computing the absolute light level for indoor visual tasks and scene design.
An application for estimating indoor heat map by implementing the transient heat equation.
Parametric analysis is conducted under varying thermal parameters for the transient heat simulation, and the resulting virtual heat maps are compared with real-world thermal images.
@@ 2.1 Indoor Light Estimation
Directly estimating indoor light from an indoor image is an ill-posed problem. Under natural illumination, the indoor spatially-varying light depends on the outdoor context, scene geometry, and material properties. Given a 2D perspective image, some studies focus on predicting panoramic HDR environment maps (Gardner et al. (2017)), lighting representation (Gardner et al. (2019)), and estimating HDR panoramas from LDR images (Gkitsas et al. (2020); LeGendre et al. (2019); Gardner et al. (2017)). Extensive research efforts have been devoted to estimating HDR from LDR images using inverse tone mapping algorithms (Rempel et al. (2007); Banterle et al. (2006, 2007); Reinhard et al. (2002)). The estimated HDR panorama is used as a global light source for high-quality relighting and inserting new virtual 3D objects into the scene. Previous studies have also explored indoor lighting editing (Li et al. (2022)), material property estimation (Yeh et al. (2022)), and recovery of spatially-varying lighting (Li et al. (2020); Garon et al. (2019); Srinivasan et al. (2020)) from images. User inputs provide additional information to define scene geometry and light location and direction (Karsch et al. (2011)). Through user-annotated light positions, Zhang et al. (2016) recover indoor illumination and room geometry from a complete RGB-D scan of a scene.
Since the input 2D perspective image is in LDR format, while the estimated HDR environment map from previous work serves as a global light source, it often fails to recover the accurate scene radiance of the real world. More importantly, earlier studies that directly estimate environment maps from images do not construct a complete light transport model. Such a model would ideally account for outdoor lighting, scene geometry, and material properties and present a complete interaction between light and the environment.
@@ 2.2 Outdoor Light Estimation
Outdoor light changes with the appearance of the sky, the surrounding context, and the time of day. The luminance and angular distribution of the sky also vary with cloud levels, times of year, and geographic locations. Perez sky model (Perez et al. (1993)) and the CIE standard sky model (CIE (2003)) are commonly used to estimate outdoor light. Due to the nuanced difference between cloud cover and sunlight intensity, generalized sky models cannot fully describe the realistic sky appearance (Darula and Kittler (2002)). Recent studies have focused on estimating the appearance of the spectral sky. A spectral sky represents chromatic changes based on time, sky turbidity, and geographic locations (Hosek and Wilkie (2012)). Furthermore, analytical sky models (Lalonde et al. (2012); Preetham et al. (1999)) approximate full-spectrum daylight under various atmospheric conditions.
In addition to generating sky model images, some studies aim to estimate outdoor lighting directly from 2D images. Learning-based method estimates 360∘ outdoor environment map from a single 2D photograph (Hold-Geoffroy et al. (2017); Lalonde and Matthews (2014)). Lalonde and Matthews (2014) developed a method to estimate light direction and intensity from a single Low Dynamic Range (LDR) image, and the generated HDR sky models are used for relighting new objects. Zhang and Lalonde (2017) recover linear HDR outdoor images from a single LDR panorama. Sunkavalli et al. (2008) recover time-varying color variation for the outdoor scene under direct sunlight and ambient skylight. The recent study by Dastjerdi et al. (2023) directly adds, removes, and edits the indoor and outdoor illumination sources in the panoramic representation. However, those studies focus on estimating outdoor environment maps for relighting new virtual objects; the outdoor light source is not used to estimate indoor light transport.
@@ 2.3 Thermal Imaging Technique
Above absolute zero temperature, indoor objects radiate infrared energy in the Long-Wave Infrared (LWIR) spectrum (wavelengths between 8 and 15 ), resulting in heat on the object’s surface. The radiometric quantity of the heat transfer can be measured by infrared thermal imaging technique (Holst (2000); Kaplan (2007); Vollmer (2020); Planinsic (2011)). Thermal imaging has gained increasing attention in computer vision research and leads to applications based on the thermal spectrum, such as object segmentation (Huo et al. (2023)), 3D estimation (Eren et al. (2009); Shin et al. (2023); Narayanan et al. (2025)), and material property estimation (Dashpute et al. (2023)).
Thermography has been extensively used to collect surface heat in building spaces. In outdoor applications, recent studies have focused on identifying thermal leakage on exterior facades (Motayyeb et al. (2023)) and integrating thermal attributes with 3D point clouds (Jarząbek-Rychard et al. (2020)). In indoor scenarios, recent research has combined 3D point clouds with 2D thermography for generating 3D building models (Adán et al. (2017); Schmoll et al. (2022)), detecting faults on interior surfaces (Garrido et al. (2018)), and segmenting targeted thermal regions on building surfaces (Adán et al. (2020)). Additionally, heat maps can be captured by a thermal-infrared camera from multiple indoor locations to analyze the heat generated on vertical wall planes (López-Fernández et al. (2017)), support 3D modeling (Adán et al. (2017)), and combine with RGB-D camera for spatial mapping (Vidas et al. (2013)).
Existing studies on indoor thermography primarily focus on capturing temporal heat maps at specific time points. There are several limitations. First, 2D thermography captures only a limited field of view. Second, capturing multiple 2D perspective images requires additional effort in data post-processing, such as image registration and alignment. Third, implementing infrared thermal cameras with additional devices like RGB-D cameras or LiDAR adds complexity and increases data acquisition costs.
@@ 3.1 Indoor Light Estimation
Using a single panorama, scene editing enables users to modify the indoor camera position, adjust lighting conditions, and change materials. Fig. 3 illustrates a reading scenario and shows the effects of indoor changes with three desk materials: glass, marble, and wood. Beyond the RGB image, the absolute light level is crucial for indoor daily activities. The virtual scene is rendered in HDR format and then converted into luminance values through per-pixel computation (Fig. 4). Since both indoor and outdoor HDR images are calibrated to accurately represent real-world scene radiance, this process ensures that the luminance value () in the virtual rendered scene is accurately estimated through physics-based rendering.
3.2 Indoor Task Lighting
In addition to luminance (scene radiance), illuminance (scene irradiance) is commonly used to evaluate the adequacy of indoor task lighting. Each indoor task has its own specific target light level to ensure optimal lighting conditions. To convert and obtain the illuminance map, we assume the scene object is Lambertian surface, where the reflectance () of the diffuse scene materials are known. The global illuminance map can be derived from displayed luminance value on HDR image(Yang et al. (2013)).
@@ 3.2 Indoor Task Lighting
The indoor task is visualized from a 180∘ hemispherical fisheye perspective. Fig. 5 showcases a virtual scene for a reading task. When the desk in the original scene is repositioned to face the wall, and a new desk lamp and ceiling light are added to illuminate the scene, the luminance levels in the virtual scenes can be estimated from the virtual HDR image. Based on the Lambertian assumption for the scene, the average illuminance level on the desk surface can be calculated. If the reading task requires a target brightness of 300 , a code-based compliance analysis computes the error map to evaluate whether the light is above or below the target level for the reading task.
4 Indoor Heat Estimation
The heat estimation requires absolute energy influx for each indoor location. As shown in Fig. 6, we rendered the existing scene in HDR format and then used per-pixel computation to convert it into the corresponding luminance map. Typically, converting illuminance to energy influx requires specialized measurement equipment. Because we focus on the indoor scene under natural light, the per-pixel energy influx can be approximated directly from the illuminance values (obtained from Equ. 1). In indoor settings, an engineering conversion suggests that 120 is approximately equivalent to 1 for sunlight (Michael et al. (2020)). As shown in Fig. 6, the energy influx, converted from the indoor luminance map, will be used as an absolute heat source to compute the indoor heat map.
@@ 4.1 Transient Heat Simulation
Where is the thermal conductivity of the material (), is the density of the material (), and is the specific heat capacity of the material at constant pressure (). The indoor scene has a wide range of building materials, and those variables can be accessed from the existing material database (IESVE (2018)).
Intrinsic attributes, such as the Laplace operator, are fundamental quantities in various physical simulations, including heat transfer. The work by Narayanan et al. (2025) estimates intrinsic 3D shapes from thermal videos, where transient heat transport follows:
The first term represents conduction, and the second term represents the sum of radiation, convection, and input heat source; other quantities are explained in Table 1.
The Equ. 3 applies to the object heated by three transfer approaches given the surrounding temperature. For indoor surfaces, heat changes not only by spatially-varying light distribution but also by heat transfer between indoor and outdoor spaces. Therefore, we include heat exchange between indoor and outdoor spaces in the heat equation and add outdoor temperature () into the heat transport equation:
@@
We estimate the indoor heat map using the transient heat equation to calculate the temperature at each pixel. In Fig. 7, the 3D floor layout is derived from a single indoor panorama. The heat simulation is conducted in 3D coordinates, with planar surfaces subdivided into vertices to compute the heat value for each point. After the simulation, the heat data in 3D coordinates is mapped onto an equirectangular projection and displayed on a panorama with the same camera position as the input panorama.
@@
Our application can estimate indoor light and heat panoramas based on new outdoor images. In Fig. 9, the existing scene is rendered under real-time outdoor lighting at 13:48. We virtually edited the sun’s position to simulate different times and estimated the corresponding panoramic heat map. In Fig. 10, we present the joint estimation of indoor light and heat properties within the scene. The existing scene is renovated with new furniture objects and then converted into the corresponding light and heat maps. Following the approach by Ji et al. (2024), the outdoor scene is initially cloudy, and direct sunlight is added to the outdoor image and used as a light source to relight the scene at different times in the afternoon (16:45, 17:45, and 18:45). In contrast to the cloudy day scenario, increased direct sunlight enters the indoor space, intensifying illumination on the table and walls while producing strong reflections on the refrigerator. This direct sunlight also raises the temperature levels on the wall and floor surfaces near the window.

### v2

# Indoor Heat Estimation from a Single Visible-Light Panorama

@@ Abstract
This paper introduces a novel image-based rendering technique for jointly estimating indoor lighting and thermal conditions from paired indoor-outdoor high dynamic range (HDR) panoramas. Our method uses the indoor panorama to estimate the 3D floor layout, while the corresponding outdoor panorama serves as an environment map to infer spatially-varying illumination and material properties. Assuming indoor surfaces are Lambertian and that all heat originates from outdoor visible light, we model the relationship between light transport and heat transfer, and perform transient heat simulations to generate indoor temperature distributions. The simulated heat maps are validated against real-world thermal images captured with an infrared camera. This approach supports photorealistic and physically informed visualization, enabling integrated light and heat estimation to advance traditional virtual home staging.
1 Introduction
Omnidirectional photography enables the capture of a single image that comprehensively represents the entire indoor scene. Recent studies [59, 29, 30, 31] have shown that a single indoor panorama can be used to estimate spatially-varying light, material properties, and 3D floor layouts, supporting applications such as indoor virtual staging. As outdoor light enters a space, it not only shapes the visual appearance of the interior, but it also introduces solar radiation that contributes to heat accumulation on indoor surfaces. Estimating the heat distribution is essential for evaluating the potential heating and cooling demands throughout the space. However, thermal assessment typically relies on infrared thermal cameras, which are expensive, prone to noise, and limited to capturing 2D perspective images with narrow field of view (FOV). These limitations hinder the ability to obtain high-quality thermal information across an entire 360∘ indoor scene. This challenge motivates a key research question: Can we directly estimate indoor thermal properties from a single visible-light panorama?
@@ 1 Introduction
Real-world lighting varies over time with spectral characteristics. To model outdoor illumination, virtual sky models [8, 42, 22] have been developed for light estimation. The global rendering method [12] provides a high dynamic range (HDR) image-based rendering setup for relighting virtual objects within realistic scene contexts. More recently, image-based rendering techniques have used captured or estimated HDR images to achieve instantaneous light estimation from images. HDR techniques capture the appearance of the entire sky, including direct sunlight and atmospheric conditions, allowing an accurate representation of the luminance distribution of the sky and the surrounding urban context [25, 26]. By using such outdoor HDR images as illumination sources, indoor environments can be rendered at the pixel level with realistic lighting effects.
@@
Intrinsic decomposition models light transport by splitting an image into reflectance (albedo) and shading (illumination) components. Intrinsic decomposition has been applied in various tasks, including image editing [6], removing interreflections [49], and separating illumination for material recoloring [7], as well as being applied to multi-view outdoor photos [13]. To construct indoor-outdoor light transport, previous studies [29, 30] used an indoor panorama and a paired outdoor hemispherical image to reconstruct the indoor appearance. In this work, we simplify the rendering setup using indoor-outdoor panoramas captured with a single camera. This method streamlines data collection and post-processing.
@@
Heat transport includes the process of heat generated by the source and its exchange between different mediums [5]. Temperature exchange within an object follows the well-known heat transfer physics, and three primary mechanisms include conduction, convection, and radiation. Recent work by Ramanagopal et al. [45] bridges visible light transport and heat transport in solids and estimates heat generation from light absorption. Previous studies on heat estimation have primarily focused on small objects under controlled artificial lighting and captured thermal images in a 2D perspective. In this work, we estimate the absolute energy influx on indoor surfaces using photometrically calibrated indoor–outdoor HDR images and compute the resulting heat generated by outdoor light on these surfaces.
In this work, using paired indoor–outdoor HDR panoramas as input, we combine physics-based and data-driven methods to estimate both light and heat properties within indoor environments. The proposed approach supports applications such as virtual home staging and immersive visualization (Fig. 1). By estimating heat from visible-light information, the method quantifies indoor thermal effects contributed by outdoor natural light. In summary, this study makes two primary technical contributions:
Visible-Light Estimation We propose a method that establishes a connection between indoor light transport and heat transfer. Using photometrically calibrated indoor–outdoor HDR panoramas, our application reconstructs radiometrically consistent indoor scenes and estimates spatially varying lighting and thermal properties for panoramic visualization (Fig. 2).
Transient Heat Estimation We implement a transient heat simulation model that quantifies the contribution of solar heat gain to indoor thermal distribution. The model accounts for conduction, convection, and radiation, and can be customized through user-defined thermal parameters. The simulation produces panoramic heat maps that change over time. The estimated heat maps are validated against thermal images captured in real environments.
@@ 2.1 Indoor Light Estimation
Directly estimating indoor light distribution from a single image is an ill-posed problem. Given a 2D perspective image, some studies focus on predicting panoramic HDR environment maps [16, 19, 35, 15], and extensive research efforts have been devoted to estimating HDR from LDR images using inverse tone mapping algorithms [47, 3, 4, 46]. The estimated HDR panorama is used as a global light source for high-quality relighting and inserting new virtual 3D objects into the scene. Previous studies have also explored indoor lighting editing [37], material property estimation [57], and recovery of spatially-varying lighting [36, 17, 51] from images.
Since the input 2D perspective image is in LDR format, and the estimated HDR environment map from previous work is used as a global light source, the resulting scene often fails to reflect accurate real-world radiance. More importantly, earlier studies that estimate environment maps directly from images do not construct a complete light transport model. An ideal model should account for outdoor lighting, scene geometry, and material properties to fully represent the interaction between light and the environment.
@@ 2.2 Outdoor Light Estimation
Outdoor light varies with sky appearance, surrounding context, and time of day. Luminance and angular distribution of the sky are further influenced by cloud cover, season, and geographic location. Perez sky model [42] and CIE standard sky model [8] are widely used to estimate outdoor illumination. However, due to subtle differences in cloud cover and sunlight intensity, generalized models cannot fully capture realistic sky appearance [9]. Recent studies have aimed to estimate spectral appearance of the sky, which accounts for chromatic variations based on time, sky turbidity, and geographic location [22]. In addition, analytical models [33, 44] have been developed to approximate full-spectrum daylight under diverse atmospheric conditions.
Some studies aim to estimate outdoor lighting directly from 2D images. Learning-based methods estimate 360∘ outdoor environment maps from a single 2D photograph [20, 34], and the resulting HDR sky models are used for relighting tasks. Zhang and Lalonde [58] recover linear HDR outdoor images from a single LDR panorama. Sunkavalli et al. [52] reconstruct time-varying color changes in outdoor scenes under direct sunlight and ambient skylight. More recently, Dastjerdi et al. [11] proposed editing indoor and outdoor illumination sources directly in a panoramic representation. However, these studies primarily focus on estimating outdoor environment maps for relighting virtual objects, without modeling indoor light transport from outdoor sources.
@@ 2.3 Thermal Imaging Technique
Above absolute zero temperature, indoor objects radiate infrared energy in the Long-Wave Infrared (LWIR) spectrum (wavelengths between 8 and 15 ), resulting in heat on the object’s surface. The radiometric quantity of the heat transfer can be measured by infrared thermal imaging technique [21, 32, 55, 43]. Thermal imaging has gained increasing attention in computer vision research and leads to applications based on the thermal spectrum, such as object segmentation [23], 3D estimation [14, 50, 41], and material property estimation [10]. Meanwhile, thermography has been widely used in building analysis, from detecting facade heat loss [40] to integrating thermal data with 3D point clouds [28]. In indoor settings, it supports 3D modeling [2, 48], fault detection [18], thermal region segmentation [1], and spatial mapping using thermal and RGB-D cameras [54, 38].
Existing studies on indoor thermography primarily focus on capturing temporal heat maps at specific time points, but several limitations remain. First, 2D thermography provides only a limited field of view. Second, acquiring multiple 2D perspective images requires additional post-processing, such as image registration and alignment. Third, integrating infrared cameras with devices like RGB-D sensors or LiDAR increases system complexity and data acquisition costs.
@@ 3.1 Indoor Light Estimation
Beyond the RGB image, the absolute light level is crucial for indoor activities. The virtual scene is rendered in HDR format and then converted into luminance values through per-pixel computation [27]. Since both indoor and outdoor HDR images are calibrated to accurately represent real-world scene radiance, this process ensures that the luminance value () in the virtual rendered scene is accurately estimated through physics-based rendering.
In addition to luminance (scene radiance), illuminance (scene irradiance) is commonly used to evaluate the adequacy of indoor task lighting. Each indoor task has its own specific target light level to ensure optimal lighting conditions. To convert and obtain the illuminance map, we assume the scene object is Lambertian surface, where the reflectance () of the diffuse scene materials are known. The global illuminance map can be derived from displayed luminance value on HDR image [56].
@@
3.2 Indoor Heat Estimation
The heat estimation requires absolute energy influx for each indoor location. We rendered the existing scene in HDR format and then used per-pixel computation to convert it into the corresponding luminance map. Typically, converting illuminance to energy influx requires specialized measurement equipment. Because we focus on the indoor scene under natural light, the per-pixel energy influx can be approximated directly from the illuminance values by using paired HDR images and a physics-based rendering process (obtained from Eq. 1). The influx () of solar energy is converted from illuminance () caused by sunlight. We apply an engineering rule of thumb that 120 equals 1 [39]. The per-pixel energy influx will be used as an absolute heat source to compute the indoor heat map.
@@ 3.3 Transient Heat Simulation
Where is the thermal conductivity of the material (), is the density of the material (), and is the specific heat capacity of the material at constant pressure (). The indoor scene has a wide range of building materials, and those variables can be accessed from the existing material database [24].
Intrinsic attributes, such as the Laplace operator, are fundamental quantities in various physical simulations, including heat transfer. The work by Narayanan et al. [41] estimates intrinsic 3D shapes from thermal videos, where transient heat transport follows:
Here, is an indicator function that is 1 if the volume is exposed to the surface and 0 otherwise. The first term represents conduction, and the second term represents the sum of radiation, convection, and input heat source; other quantities are explained in Table 1.
The Eq. 3 applies to the object heated by three transfer approaches given the surrounding temperature. For our indoor heat estimation, we provide per-pixel energy influx as the heat source and apply the remaining thermal properties using known values. Besides, for indoor surfaces, heat changes not only by spatially-varying light distribution but also by heat transfer between indoor and outdoor spaces. Therefore, we include heat exchange between indoor and outdoor spaces in the heat equation and add outdoor temperature () to the heat transport equation:
@@
We estimate the indoor heat map using the transient heat equation (Eq. 4) to calculate the temperature at each pixel. In Fig. 3, the 3D floor layout is derived from a single indoor panorama. The heat simulation is conducted in 3D coordinates, with planar surfaces subdivided into vertices to compute the heat value for each point. After the simulation, the heat data in 3D coordinates is mapped onto an equirectangular representation and displayed from the same camera position as the input panorama.
@@
Our application can estimate indoor light and heat panoramas based on new outdoor images. In Fig. 5, we present the joint estimation of the indoor light and heat properties for the scene. The existing scene is renovated with new furniture objects and then converted into the corresponding light and heat maps. Following the approach of Ji et al. [30], the outdoor scene is initially cloudy, and direct sunlight is added to the outdoor image and used as a light source to relight the scene at different times in the afternoon (16:45, 17:45 and 18:45). Compared to the cloudy scenario, direct sunlight increases indoor light levels on the table and walls, leading strong reflections on the refrigerator, while also increasing temperatures on wall surfaces near the window.
The transient heat simulation can be performed with different surface materials [24]. As shown in Fig. 6, we first used plaster (dense) as the surface material for the scene, and the resulting estimated heat map served as the baseline. We then replaced the surface material with glass and wood, generating the corresponding heat maps for the same scene. Compared with plaster (dense), both glass and wood produced lower surface temperatures.
An accurate 3D floor layout is essential for precise heat estimation. We focus on the error analysis of the heat simulation in different floor layouts (Fig. 7). Layout 1 is considered as the correct floor layout, while Layouts 2 and 3 feature slight shifts in the floor boundaries. The heat maps for Layouts 2 and 3 reveal variations in indoor temperature values due to inaccuracies in the room geometry. In the error map, the heat maps of Layouts 2 and 3 are compared with the ground-truth (GT) heat map of Layout 1. Layout 2 shows lower temperature estimates at the upper and lower corners of the wall, while Layout 3 exhibits lower temperatures near the wall adjacent to the left window. The 3D layout estimation assumes a Manhattan layout, which is effective for regular indoor geometries. Curved walls and tilted roofs are simplified into a cubic model.